### v1

# FMDA-OT: Federated Multi-source Domain Adaptation Through Optimal Transport

Abstract
Multi-source Domain Adaptation (MDA) aims to adapt models trained on multiple labeled source domains to an unlabeled target domain. In this paper, we introduce our approach as a collaborative MDA framework, which comprises two adaptation phases. Firstly, we conduct domain adaptation for each source individually with the target, utilizing optimal transport. Then, in the second phase, which constitutes the final part of the framework, we design the architecture of centralized federated learning to collaborate the N models representing the N sources. This architecture offers the advantage of using the sources without accessing their data, thus resolving data privacy issues inherent in domain adaptation. Additionally, during this phase, the server guides and fine-tunes the adaptation using a small number of pseudo-labeled samples available in the target domain, referred to as the target validation subset of the dataset.
Keywords Federated Learning  $\cdot$ Multi-source Domain Adaptation  $\cdot$ Optimal Transport  $\cdot$ Data Science  $\cdot$ Machine Learning.
1 Introduction
Supervised learning is a pillar of machine learning, thus this supervision requires labels, which is most of the time an expensive task. While transfer learning and generalisation can be seen as solutions, but due to domain shift, in other words, divergence between the domains (i.e. the difference between probability distributions of two domains), the task model trained on labelled data from domain A (source) and tested on unlabelled data from domain B (target) will have a decay in the performance in comparison with training and testing on the same domain B, which is out capacity in the absence of labels. To address these challenges, domain adaptation approaches are employed. Within these approaches, some can manage adaptation between a single source and a single target, referred to as single-source domain adaptation (SDA) methods, while others can handle adaptation between multiple sources and a single target, known as Multi-source domain adaptation (MDA) methods.
Several works have been proposed in MDA, [Mansour et al., 2021] defended a theoretical analysis considering the assumption says that the target is a convex combination of source domains. [Zhao et al., 2019] Proposed new error bounds for both regression and classification tasks. There are other works proposed on specific algorithms. [Zhao et al., 2020] Combines diverse classifiers by a novel weighting technique. [Hoffman et al., 2012] Introduced a domain transform framework that finds the latent space of the domains by using a clustering approach. These researches generally, either estimate a mixed distribution or combine with a specific approach multiple single-source models.
Collaborative learning paradigm generally and Federated Learning (FL) particularly has attracted the attention of the academia and industry communities. The term Federated Learning was introduced by [McMahan et al., 2017]. This new machine learning approach permits to train a model on different data from different sources that can be mobile devices or even organizations. Our implemented version of federated learning will be used in our domain adaption approach, where each source domain will be considered and represented as a data source in the federated learning paradigm. This approach will consider and open the door for data privacy in domain adaptation where we do not have the access to the source domains data.
Unsupervised domain adaptation methods relying on optimal transport have gained attention due to the recent success of optimal transport in diverse machine learning problems [Laclau et al., 2017] [Ben-Bouazza et al., 2019] [Ben-Bouazza et al., 2022]. These domain adaptation models have increasingly captured interest thanks to their ability to explicitly discover and learn the lowest transformation cost from the source to the target domains. This capability enables them to effectively reduce the discrepancy between the two domains.
Addressing this limitations we propose our approach which combines between optimal transport and federated learning to train a model on multiple source domains and have better performance on the target domain. This paper will be organised in the rest as: Section 2 offers the required preliminary and preparatory knowledge beside notations for unsupervised multi-source domain adaptation, as well as optimal transport theory for domain adaptation. Section 3 has been devoted to present the proposed approach for MDA, which is named Fedetared Multi-source Domain Adaptation through Optimal Transport (FMDA-OT). Section 4 demonstrates the performed comparative study with state-of-the-art methods on two benchmark datasets for MDA. To conclude in section 5.
2 Fundamental background of the proposed approach
2.1 Unsupervised Multi-Source Domain Adaptation
Let $\mathcal{X}=\mathbb{R}^{d}$ be an input space and $\mathcal{Y}=\{c_{1},...,c_{k}\}$ a set of label space consisting of $k$ classes. Let’s consider $\mathcal{S}_{i}\,\forall i\in\{1,...,N\}$, $N$ distinct probability distributions over the product space $\mathcal{X}\times\mathcal{Y}$ called the source domains, and $\mathcal{T}$ a probability distribution over the product space $\mathcal{X}\times\mathcal{Y}$ called the target domain.
Unsupervised Multi-source domain adaptation (UMDA) is a transductive learning problem, therefore, both labelled data from the source domains and unlabelled data from the target domain are assumed to be accessible throughout the training phase [Redko et al., 2019].
More formally, for each source, we have access to a set $S_{i}=\{(x_{j}^{i},y_{j}^{i})\}_{j=1}^{n_{i}}$ of $n_{i}$ independent and identically distributed (i.i.d.) labeled samples of the $i^{th}$ source pulled from the joint distribution $\mathcal{S}_{i}$, and a set $T=\{x_{j}\}_{j=1}^{m}$ of $m$ i.i.d. unlabeled target samples pulled from the marginal distribution $\mathcal{T}_{\mathcal{X}}$, of the joint distribution $\mathcal{T}$ over $\mathcal{X}$, properly:
[…]
The objective of unsupervised multi-source domain adaptation is to build a robust classifier $\eta:\mathcal{X}\to\mathcal{Y}$ with a slight target risk:
$\mathcal{R}_{\mathcal{T}}(\eta)=\underset{(x,y)\sim\mathcal{T}}{\mathbb{P}}(\eta(x)\neq y)$,
Under the assumption of domain shift not only between the source and target domains but also among the source domains themselves:
[…]
2.2 Optimal Transport
Let’s consider two probability measures $\mu$ and $\nu$ over spaces $\mathcal{X}$ and $\mathcal{Y}$ respectively, and a positive cost function $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{+}$. The aim of optimal transport is to minimize the total cost of transporting the measure $\mu$ into $\nu$.
In the discrete formulation of optimal transport, when $\mu$ and $\nu$ are only available through finite samples $X=\{x_{1},x_{2},...,x_{n}\}$ and $Y=\{y_{1},y_{2},...,y_{m}\}$, they can be considered as empirical distributions:
$\mu=\sum_{i=1}^{n}\alpha_{i}\delta_{x_{i}}\,\,\text{and}\,\,\nu=\sum_{j=1}^{m}\beta_{j}\delta_{y_{j}}\vspace{3mm}$ (1)
where $\alpha=(\alpha_{1},...,\alpha_{n})$ and $\beta=(\beta_{1},...,\beta_{m})$ are vectors in the probability simplex $\sum_{n}$ and $\sum_{m}$ respectively. In this case, the cost function only needs to be specified for every pair $(x_{i},y_{j})_{i,j}$, so it can be compactly represented by a cost matrix $C\in\mathcal{M}_{n\times m}(\mathbb{R}^{+})$.
The formulation of Monge aims to get a measurable transport map $\texttt{T}:\mathcal{X}\to\mathcal{Y}$ that assigns each point $x_{i}$ to a single point $y_{j}$, and which pushes the mass of $\mu$ toward that of $\nu$ (i.e., $\texttt{T}\#\mu=\nu$), while minimizing the total cost of transportation:
[…]
where ##\## denotes the pushforward operator.
Since the problem of Monge may not admit a solution, a convex relaxation was suggested by Kantorovich that considers instead of the transport map T a soft multi-valued assignment defined in terms of probabilistic couplings $\gamma\in\mathcal{M}_{n\times m}(\mathbb{R}^{+})$ whose marginals recover $\mu$ and $\nu$. Formally, Kantorovich’s formulation seeks $\gamma$ in the transportation polytope:
$U(\alpha,\beta)=\{\gamma\in\mathcal{M}_{n\times m}(\mathbb{R}^{+})\,|\,\gamma\mathds{1}_{m}=\alpha\,\text{and}\,\gamma^{\mathrm{T}}\mathds{1}_{n}=\beta\}$ (3)
that solves:
$(\mathcal{MK})\,\,\,\underset{\gamma\in U(\alpha,\beta)}{\min}\langle{\gamma},{C}\rangle_{F},$ (4)
where .,.$\langle.,.\rangle_{F}$ is Frobenius inner product of matrices.
The discrete Kantorovich problem is a linear program and can therefore be solved exactly in $\mathcal{O}(r^{3}log(r))$, where $r=\max(n,m)$ using simplex algorithm and interior point methods, that is a heavy calculation price. In practice, An entropy smoothing variant [Cuturi, 2013] leads to much more efficient algorithms:
$(\mathcal{MK}^{\,\varepsilon})\,\,\,\,\,\,\underset{\gamma\in U(\alpha,\beta)}{\min}\langle{\gamma},{C}\rangle_{F}-\varepsilon\mathcal{H}(\gamma),\vspace{2mm}$ (5)
where $\mathcal{H}(\gamma)=-\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}(\log(\gamma_{ij})-1)$ is the entropy of $\gamma$. The solution to this strictly convex optimization problem is of the form $\gamma^{*}=diag(u)Kdiag(v)$, with $K=\exp(\frac{-C}{\varepsilon})$, and can be obtained efficiently via the Sinkhorn-Knopp algorithm, an iterative matrix scaling method [Peyré et al., 2019].
2.3 Federated Learning
Federated learning enables us to jointly train a machine learning model through the collaboration of multiple parties (i.e. devices, organisations …etc.) without exchanging the local data. Generally federated learning approaches can be classified into either centralised or decentralised paradigm. Many federated learning systems today support neural network training. Google proposed a scalable production system that authorizes the training of deep neural networks by hundreds of millions of devices [Bonawitz et al., 2019]. Centralised FL can be represented as a set of N models for a set of N clients (i.e data sources) and a server model that centralises the learning process. Figure 1 shows the centralised FL paradigm.
Figure 1: Centralised Federated Learning
The training process in centralized federated learning (FL) can be viewed as consisting of four repetitive phases. Firstly, each client model is trained on its local private data. Then, as a second phase, each client sends its model (i.e., weights) to the server following the FedAvg algorithm [McMahan et al., 2017]. Various algorithms have been proposed to handle the exchange of information between models, one of which is the FedSGD algorithm [McMahan et al., 2017]. This approach allows for the transmission of gradients rather than weights, which is considered more effective but requires increased communication, potentially straining the distributed FL system. The third phase involves the server averaging the received weights from the N clients. This average is weighted by the number of samples in each client’s data, normalized by the sum of the counts of all samples across all clients, as proposed by Google [McMahan et al., 2017]. Finally, in the fourth phase, the clients receive the new weights and update their models.
The decentralized or Peer-to-Peer federated learning paradigm follows the same principles as centralized FL, except it doesn’t require a trusted server to centralize the learning process. Figure 2 illustrates the overall repetitive process of training N client models.
Figure 2: Decentralised Federated Learning
Initially, each client trains its model on its local private data. Then, in the second and third phases, it requests the weights or gradients from other clients to update its model with its local data and the received exchangeables (i.e., weights or gradients). Many algorithms have been proposed in this paradigm; BrainTorrent [Roy et al., 2019] is a selective algorithm in which clients communicate with each other and update their models using only the exchangeables from clients who have updates to send. This process continues until convergence. It’s evident that the order of operations is crucial in decentralized federated learning algorithms and significantly impacts the final training results. The outcome may differ if, for instance, client 1 uses the weights or gradients from client 2 before client 2 does. This challenge emphasizes the importance of collaboration order. simFL [Li et al., 2020] is a decentralization of federated learning that adopts the centralized paradigm in a decentralized manner. The proposal involves selecting one client as the server in each iteration and treating the others as clients. The model is then updated based on the chosen client’s weights and the received ones, before being sent back to the rest. However, simFL may encounter security issues or instability due to the variable server.
3 Our Approach
3.1 Overall framework of FMDA-OT
Figure 3 illustrates the overall framework of our proposition, which can be divided into two main parts of the adaptation. Initially, each client from the source domains receives a small amount of target domain data, which will be pseudo-labeled for reasons discussed later in the paper. In this phase, we apply the optimal transport to project each source domain data into the target domain space using the Sinkhorn algorithm with L1L2 class regularization. This process aims to find a new representation of the source domain data that closely resembles the representation of the target domain, thereby reducing the domain shift phenomenon between them. This new representation is referred to as the transported data of the source domain.
Figure 3: Overall Framework of FMDA-OT
Optimal transport generates new representations of the source domain data, which may not always be the most discriminative representation. In some cases, it could be inferior to the original representation of the source domain data, possibly due to factors such as data imbalance. To address this issue, we utilize pseudo-labeled data from the target domain to assess the quality of the new representation. This involves testing and comparing the accuracy of models trained on both the transported and original data. If an improvement is observed when testing on the transported data, the client may opt to use this data for further step of the adaptation. Otherwise, the original representation is retained as the preferred choice. The subsequent step involves the intervention of centralized federated learning, with each client acting as a source domain. In our approach, we employ the FedAvg algorithm for the learning process. Here, each client sends its own weights after the local training step, and the server aggregates these weights, taking into account coefficients assigned to each client.
Google, in its proposal [McMahan et al., 2017], defines the weight aggregation that takes place on the server as a weighted sum of all the clients’ weights. The coefficient assigned to each client is directly proportional to the number of samples in that client. Formally:
$W_{s}=\sum_{i=1}^{N}\frac{n_{i}}{S_{N}}W_{i}$ (6)
with :
$S_{N}=\sum_{i=1}^{N}n_{i}$ (7)
While it is logical and sensible to weight the sum of the weights based on the number of samples provided by each client, this approach appears to work better in scenarios where all clients have the same probability distribution (i.e., the same domain). In unsupervised Multi-source Domain Adaptation (MDA), where different domains are involved, a client with fewer samples may contribute more effectively to improving the final model than another client with more data. Failing to account for this domain shift and weighting solely based on the number of samples can lead to less reliable results. To address this issue, we propose weighting all clients by defining coefficients that are directly proportional to the performance of each client model when tested on the pseudo-labeled portion of target domain data.
Our proposition takes into account the paramount importance of privacy, which is a necessity across all industries today. It is often challenging to train a model without direct access to the data, particularly in the context of domain adaptation and phenomena such as domain shift, where data from multiple sources may be required to leverage information for building a robust model while still adhering to privacy regulations. To enhance domain adaptation within the constraints of privacy, we propose FMDA-OT. The following Table 1 summarizes the advantages introduced by FMDA-OT.
Table 1: Adaptation settings comparison
Settings
Method Name Unlabelled Target Data No Source Data Access Source Model Multiple Domains
HTL[Singh et al., 2018]
UDA[Hoffman et al., 2018]
MSDA[Peng et al., 2019]
U-HTL[Liang et al., 2020]
STEM[Nguyen et al., 2021b]
MOST[Nguyen et al., 2021a]
FMDA-OT (Ours)
3.2 Pseudo labelling the validation data of target domain
The significance of this validation proportion lies in its ability to guide adaptation, thereby reducing domain shift by maintaining a dynamic training process. In our scalable proposition, this validation proportion can be labeled by experts if available, which can lead to improved results. Alternatively, we propose a pseudo-labeling approach to handle full unsupervised training for the target domain model. The HOT [El Hamri et al., 2022] approach offers a scalable solution that can be applied in our scenario. Initially, we cluster the target domain data on the server using one of the proposed clustering algorithms. In our experiments, we utilized spectral clustering, as it represents the state-of-the-art in clustering algorithms. Once the clusters are obtained and random labels are assigned, this pseudo-labeled validation data is distributed to each client. Here, we apply hierarchical optimal transport (HOT), which calculates the correspondence between each source domain label and the random labels assigned to this data during clustering. Figure 4 illustrates the HOT correspondence matrix.
Figure 4: HOT correspondence Matrix
Not necessarily do we obtain identical results from one client to another. For this reason, if the number of clients is three or more, we apply a majority vote. In cases where we have only two clients, we consider the correspondence of the closest source domain to the target according to the Wasserstein distance.
3.3 On-clients Testing And Server Weights Aggregation
The validation part of the target data serves as a guide for adaptation in both phases. During the optimal transport phase, this part determines whether accuracy improves when using the new representation obtained by the optimal transport transformation. If not, the original data representation is retained. This testing occurs across all clients with all sources. Each client receives this pseudo-labeled data and conducts an internal test to determine which data to use in the next phase of adaptation, which is federated learning.
In the federated learning training phase, the server utilizes this part. Upon receiving weights from all clients, the server tests the received model from each client on this validation part and records the accuracy. This process is repeated for all client weights. The accuracy of all clients in the federated learning is then normalized by their sum and used as weights for aggregating all incoming weights. This approach maintains highly dynamic adaptation, adjusting the model with changes occurring in each client model during the learning process. It also provides a robust metric for the importance of each source to the target.
Besides privacy, federated learning’s collaborative aspect is a significant advantage. Models share weights, addressing each model’s shortcomings. Through collaboration, we can harness the full potential of data and knowledge without sharing actual data or compromising privacy.
The guarantee of this pseudo-labeling is directly related to clustering quality. Higher cluster purity leads to better label assignments. Hence, we used spectral clustering in experiments. HOT [El Hamri et al., 2022] effectively projects random labels from clustering to real ones. This pseudo-labeling keeps the training process entirely unsupervised, yet flexible enough to replace the validation part with expert-labeled data for better weight fine-tuning. In cases of pure clusters, pseudo labels match real ones perfectly.
4 Experiments
4.1 Datasets
We conducted our experiments using established and widely recognized benchmark datasets, specifically Office-Caltech-10 and VLSC.
VLSC: Dataset consists of images from four different domains: Caltech101, VOC2007, SUN09, and LabelMe. Each domain contains five classes: birds, dogs, chairs, cars, and humans.
Office-Caltech10: Collection consists of images representing four different domains: Amazon, DSLR, Webcam, and Caltech. Each domain comprises 10 categories (i.e., classes) commonly encountered in daily life, household, and office settings, such as computer, bag, car, bike, etc.
4.2 Results and discussion
In all experiments, we test each model on each target multiple times, and then rely on the average. Currently, the number of tests is 1000.
Table 2: FMDA-OT results on VLSC
Target
Method Name
[…]
AVG
CMSD[Liu et al., 2021]
37.30
52.74
34.96
31.02
39.00
DS[Liu et al., 2021]
39.52
51.44
41.87
36.44
42.31
TCA+CMSD[Liu et al., 2021]
65.08
54.63
56.69
80.84
64.31
TCA+WAF[Liu et al., 2021]
66.67
54.81
56.40
80.53
64.60
TCA+WDS[Liu et al., 2021]
65.68
56.71
59.71
81.22
65.83
TCA+WDSC[Liu et al., 2021]
65.82
59.69
61.76
81.00
67.05
FMDA-OT (Ours)
69.0
61.0
67.0
96.0
73.25
Table 3: FMDA-OT results on Caltech-Office-10
Target
Method Name
[…]
AVG
ResNet-101[He et al., 2016]
85.4
88.7
99.1
98.2
92.9
DAN[Long et al., 2019]
89.2
91.6
99.5
99.1
94.8
DCTN[Xu et al., 2018]
90.2
92.7
99.4
99.0
95.3
MCD[Saito et al., 2017]
91.5
92.1
99.5
99.1
95.6
$M^{3}SDA$[Peng et al., 2019]
92.2
94.5
99.5
99.2
96.4
FMDA-OT (Ours)
92.0
95.1
99.4
99.2
96.5
The entire system comprises several parameters to adjust, starting with the number of neighbors in spectral clustering, which needs adjustment to enhance clustering quality. We can rely on metrics like the Silhouette index or other internal indices for this evaluation. The regularization parameters in optimal transport, either the Entropy-Regularization or the Class-Regularization, must be adjusted for each source (recommended), or they can have the same values across all sources. However, certain regularization values may lead to mathematical issues, such as division by zero. For our experiments, we provide recommendations regarding the regularization values for each source to enhance the new representation obtained by optimal transport. These parameters can be adjusted locally in each source using the pseudo-labeled data sent from the server (test-driven approach). Generally, it is preferable to minimize the Entropy-Regularization and maximize the Class-Regularization without encountering mathematical errors. In each source, minimizing the Entropy-Regularization below a certain threshold may result in mathematical operational errors. Similarly, maximizing the gap between the Entropy-Regularization and the Class-Regularization increases the likelihood of mathematical operational errors. These mathematical errors are beyond our expertise, as they are inherent to the optimal transport algorithm.
In order to assess the performance of our approaches, we use the Friedman test and Nemenyi test recommended in [Friedman, 1937] using the $Autorank$ package [Herbold, 2020]. The Friedman test is conducted to test the null-hypothesis that all approaches are equivalent in respect of accuracy. If the null hypothesis is rejected, then the Nemenyi test will be performed. In addition, if the average ranks of two approaches differ by at least the critical difference(CD), then it can be concluded that their performances are significantly different. In the Friedman test, we set the significant level ${\alpha=0.05}.$ The figure 5 shows a critical diagram representing a projection of average ranks approaches on enumerated axis. The approaches are ordered from left (the worst) to right (the best) and a thick line which connects the approaches where the average ranks not significantly different (for the level of 5% significance). As shown in figure 6, FMDA-OT achieves significant improvement over the other proposed techniques since during federated learning phase it is stable and the process stops the collaboration for some views when their local quality stars to decrease, which prevents common issue of federaed approaches. Compared to the most cited approaches of the state of the art, the positive impact of using the federated learning based on this theory is:
The statistical analysis was conducted for 7 approaches with 5 paired samples. The family-wise significance level of the tests is alpha=0.050. We rejected the null hypothesis that the population is normal for the approaches CMSD (p=0.000), DS (p=0.000), TCA+CMSD (p=0.000), TCA+WAF (p=0.000), TCA+WDS (p=0.000), TCA+WDSC (p=0.000), and FMDA-OT (p=0.000). Therefore, we assume that not all approaches are normal. Because we have more than two approaches and some of them are not normal, we use the non-parametric Friedman test as omnibus test to determine if there are any significant differences between the median values of the approaches. We use the post-hoc Nemenyi test to infer which differences are significant.
Figure 5: Friedman and Nemenyi test for comparing multiple approaches over Office-Caltech10 data sets: Approaches are ordered from right (the best) to left (the worst)
MR MED MAD CI $\gamma$ Magnitude
ResNet-101 6.000 92.900 5.300 [88.700, 98.200] 0.000 negligible
DAN 4.100 94.800 4.300 [91.600, 99.100] -0.266 small
DCTN 4.100 95.300 3.700 [92.700, 99.000] -0.354 small
MCD 3.100 95.600 3.500 [92.100, 99.100] -0.405 small
FMDA-OT 2.000 96.500 2.700 [95.100, 99.200] -0.577 medium
M3SDA 1.700 96.400 2.800 [94.500, 99.200] -0.557 medium
Table 4: Summary of methods
Figure 6: Friedman and Nemenyi test for comparing multiple approaches over VLSC data sets: Approaches are ordered from right (the best) to left (the worst)
MR MED MAD CI $\gamma$ Magnitude
CMSD 6.800 37.300 2.340 [34.960, 39.010] 0.000 negligible
DS 6.200 41.870 2.350 [39.520, 42.320] -1.314 large
TCA+CMSD 4.600 64.310 7.620 [56.690, 65.080] -3.232 large
TCA+WAF 4.000 64.600 8.200 [56.400, 66.670] -3.054 large
TCA+WDS 3.000 65.680 5.970 [59.710, 65.830] -4.222 large
TCA+WDSC 2.400 65.820 4.060 [61.760, 67.050] -5.805 large
FMDA-OT 1.000 69.000 4.250 [67.000, 73.250] -6.233 large
Table 5: Summary of methods
We report the median (MD), the median absolute deviation (MAD) and the mean rank (MR) among all approaches over the samples. Differences between approaches are significant, if the difference of the mean rank is greater than the critical distance CD=1.274 of the Nemenyi test. We reject the null hypothesis (p=0.000) of the Friedman test that there is no difference in the central tendency of the approaches CMSD (MD=37.300+-2.025, MAD=2.340, MR=6.800), DS (MD=41.870+-1.400, MAD=2.350, MR=6.200), TCA+CMSD (MD=64.310+-4.195, MAD=7.620, MR=4.600), TCA+WAF (MD=64.600+-5.135, MAD=8.200, MR=4.000), TCA+WDS (MD=65.680+-3.060, MAD=5.970, MR=3.000), TCA+WDSC (MD=65.820+-2.645, MAD=4.060, MR=2.400), and FMDA-OT (MD=69.000+-3.125, MAD=4.250, MR=1.000). Therefore, we assume that there is a statistically significant difference between the median values of the approaches. Based on the post-hoc Nemenyi test, we assume that there are no significant differences within the following groups: CMSD and DS; TCA+CMSD and TCA+WAF; TCA+WAF and TCA+WDS; TCA+WDS and TCA+WDSC. All other differences are significant.
(a) VLSC datasets
(b) Office-Caltech10 datasets
Figure 7: Sensitivity Box-Whiskers plots for the all approaches
Sensitivity Box-Whiskers plots (figure 7) represents a synthesis of the scores into five crucial pieces of information identifiable at a glance: position measurement, dispersion, asymmetry and length of Whiskers. The position measurement is characterised by the dividing line on the median (as well as the middle of the box). Dispersion is defined by the length of the Box-Whiskers (as well as the distance between the ends of the Whiskers and the gap). Asymmetry is defined as the deviation of the median line from the centre of the Box-Whiskers from the length of the box (as well as by the length of the upper Whiskers from the length of the lower Whiskers, and by the number of scores on each side). The length of the Whiskers is the distance between the ends of the Whiskers in relation to the length of the Box-Whiskers (and the number of scores specifically marked). These graphs show the same overall performance behaviour observed in the case of VLSC and Office-Caltech10 datasets. They show a clear improvement as a result of the the proposed approach. This improvement is observed for all datasets used.
Conclusion & Future Direction
In our work, we propose an architecture for privacy preservation and domain adaptation in the field of multi-source domain adaptation, allowing for model training in an unsupervised manner without access to the data sources. This approach is essential in today’s landscape to uphold client privacy across various industries. The architecture consists of two adaptation phases: firstly, employing optimal transport selectively for the sources, and secondly, the server of Federated Learning (FL) guiding the adaptation by defining the importance of each source. Our approach demonstrates its effectiveness on benchmark datasets. There are several exciting directions to explore for further development and privacy preservation. Firstly, we believe that our architecture’s performance can be enhanced by developing or utilizing more sophisticated FL systems. Secondly, the new representation obtained by optimal transport in each source can be replaced by alternative representations from other algorithms that may offer greater discrimination.